The perimeter if a square is 72 inches. Find the measure of the diagonal
Found 2 solutions by richwmiller, rfer:
Answer by richwmiller(17219)

What is the formula for the perimeter of a square? 4s=72
s=18
What is the formula for the diagonal of a square?
2*18^2=d^2

Answer by rfer(16322)

72/4=18
18^2+18^2=c^2
324+324=648
c=25.5 in